OpenStudy (anonymous):

Find an exact value for sin(17pi/12)

OpenStudy (jdoe0001):

look at your Unit Circle I'd think

OpenStudy (jdoe0001):

\(\bf \cfrac{17}{12}\implies \cfrac{12}{12}+\cfrac{5}{12}\implies 1+\cfrac{5}{12}\qquad thus \\ \quad \\ \cfrac{17\pi}{12}\implies \cfrac{12\pi}{12}+\cfrac{5\pi}{12}\implies \pi+\cfrac{5\pi}{12}\)

OpenStudy (solomonzelman):

\(\color{blue}{ π=180° }\) so, \(\color{blue}{ sin( 17π/12) }\) would be same as \(\color{blue}{ sin(225°) }\) To find \(\color{blue}{ sin(225°) }\) use \(\color{blue}{ sin(180+75) }\) knowing that \(\color{red}{ sin180=0 }\) and \(\color{red}{ sin75=\frac{3\sqrt{2} }{4} }\)

OpenStudy (solomonzelman):

sin(A+B)=sin A cos B + cos A sin B

OpenStudy (anonymous):

thanks guys!

OpenStudy (solomonzelman):

You welcome !